A. Consumer Surplus from Drinking Beer Jasmine and Leonard are EC students who regularly get together at a bar for beer Friday after class. The table shows their Friday marginal benefit (MB) schedules for beer, with marginal benefit measured as number of dollars per beer. The price of a beer is $5. Jasmine and Leonard each buy two beers.e Beere First» Second< Third Fourth* Marginal Benefit e Jasmine Leonarde 20 15 13e A1. Compute Jasmines consumer surplus. Make sure you state exactly how consumer surplus is measured-not just a number, a number of something. In purchasing two beers, is Jasmine behaving rationally or irrationally? Explain your answer. A2. Compute Leonards consumer surplus. Make sure you state exactly how consumer surplus is measured-not just a number, a number of something. In purchasing two beers, is Leonard behaving rationally or irrationally? Explain your answer.d
Answer #1

Jasmine and Leonard would consume beer as long as the Marginal Benefit from the additional beer is greater than or equal to the price paid per beer.

Analyzing in this manner, Jasmine would consume 2 bottles of beer because the marginal benefit from 3rd beer is 4 which is less than the price of beer P = $5

Similarly Leonard should buy only one beer as the marginal benefit from 2nd beer is less than the price paid. But we are given that leonard buys two beers.

Ans A1.) Jasmine's consumer surplus (CS) is Calculated as (Marginal benefit - price paid)

from 1st beer, CS = 20 - 5 = 15

from 2nd beer, CS = 15 -5 = 10

therefore jasmine's total consumer surplus = 15 + 10 = 25

Jasmine is behaving rationally by consuming two beers because her marginal benefit is greater than price for both units.

Ans A2.)Leonard's consumer surplus (CS) is Calculated as (Marginal benefit - price paid)

From 1st beer, CS = 13 - 5 = 8

From 2nd beer, CS = 3 - 5 = -2

therefore leonard's total consumer surplus = 8-2 = 6

Leonard is not behaving rationally as he should only buy 1 beer because for the 2nd beer,his marginal benefit is less than the price.